OpenStudy (anonymous):

calc 3 help here is the problem: i know the answer, i just have no idea how to get it

OpenStudy (anonymous):

http://imgur.com/RE1JW

OpenStudy (anonymous):

well why do you need the explanation if you know the answer

OpenStudy (experimentx):

under transformation you would normally multiply by http://en.wikipedia.org/wiki/Jacobian_matrix_and_determinant

OpenStudy (experimentx):

it's explained here in detail http://www.youtube.com/watch?v=UZb9hZIAvL4

OpenStudy (experimentx):

say x = u^2 + v and y=2v the area would be \[\begin{vmatrix} {\partial (u^2 + v) \over \partial u} & {\partial (u^2 + v) \over \partial v}\\ {\partial (2v) \over \partial u} & {\partial (2v) \over \partial v}\\ \end{vmatrix} \times \text{ area of }[1,0] \times [0,1] \]

OpenStudy (anonymous):

awesome, thanks for the help